# A metrological framework for uncertainty evaluation in machine learning classification models

### v2

@@ 4.2.4 Numerical examples
We analyse the output predictive probabilities of the AF detection classifier using the same procedure as described in Section 4.1.4 for LC classification. The confusion matrix and box and whisker plots are given in Figures 6 and 7 respectively. The classification performance metrics are: classification loss = 0.209; EXE = 0.874; EBS = 0.622. Since the loss and other classification metrics are larger than in the case of LC classification, the median and IQR values in the uncertainty statistics are also larger. However, we see a similar pattern as the other case study in that the statistics which explicitly use the modal probability $p_{m}$ have the smallest median and IQR, and that the information entropy is the most sensitive to changes in the PMF.
@@
Figure 7: Box and whisker plot of uncertainty statistics for AF classifier.
@@ 5 Conclusion
In this paper, we have presented a metrological framework for evaluating uncertainty within machine learning classification models, motivated by the increasing use of classification models in a wide range of applications. We have identified and listed a number of sources of uncertainty within ML classification models which contribute to the output predictive uncertainty of the nominal property or variable, namely training data uncertainty, class assignment uncertainty, uncertainty about choice of model, model parameter uncertainty, and input data uncertainty, only some of which are currently addressed in the GUM suite of documents.
The work described in this paper has focused primarily on predictive uncertainty of ML classification models, with a specific focus on probabilistic ML classification models, i.e. models that return class probabilities for the output nominal variable. This focus included the evaluation of uncertainty for such nominal variables, which is not currently considered within the VIM and GUM. We have demonstrated that the PMF provides a complete characterisation of the uncertainty associated with nominal properties, and thus the predictive uncertainty in probabilistic ML classifiers. We have also motivated the use of the PMF of a nominal variable, as opposed to a summary statistic, in propagating such uncertainty within a chain of measurement models, representing either a metrological traceability chain or a processing chain for information products. We then provided example approaches using analytic or sampling methods of how to perform such propagation using the PMF.
We have reviewed various statistics that have been proposed within the literature associated with expressing the predictive uncertainty of probabilistic ML classifiers. We provided a comparison between these statistics in terms of their fundamental properties, and applied them in the context of two case studies that rely upon a metrological analysis, namely land cover classification, and atrial fibrillation detection. We found that for both case studies, the information entropy, which is the most common statistic for uncertainty within the ML community, is the most sensitive to variations in the predictive probability distributions. We have also illustrated how sources of uncertainty are addressed within the two case studies using different methodologies dependent on the classification modelling procedure.
The framework we have outlined builds upon and enhances previous work in uncertainty evaluation for nominal properties. The framework enables an extension of the VIM and GUM to uncertainty evaluation for nominal properties, and provides methods for handling the sources of uncertainty within ML classification, underpinning their use within metrological applications.
@@ Acknowledgements
The original data used for LC classification study can be found at [Morton2020a, Morton2020b, Morton2020c, Morton2020d]. More details on how this data was processed for the LC classification study are provided in [bilson2025uncertaintyawarebayesianmachinelearning]. The original data for the case study on AF detection (DeepBeat) can be found at [ashley2020deepbeat] and is described in [torres2020multi]. The custom data split used was described in [bench2024towards].
[…]

### v3

@@ 4.2.4 Numerical examples
We analyse the output predictive probabilities of the AF detection classifier using the same procedure as described in Section 4.1.4 for LC classification. The confusion matrix and summary statistics are given in Figure 5 and Table 6 respectively.
@@
The classification performance metrics are: classification loss = 0.209, EXE = 0.874, and EBS = 0.622. Since the loss and other classification metrics are larger than in the case of LC classification, the uncertainty statistics in Table 6 are also larger than those in Tables 4 and 5.
[…]
WVR $6.98$ $2.83$ $5.87$ $3.41$
UVR $4.66$ $1.88$ $3.92$ $2.28$
SDM $6.98$ $2.83$ $5.87$ $3.41$
$H$ $21.8$ $4.01$ $8.71$ $4.10$
$H^{*}$ $16.3$ $3.76$ $8.30$ $4.02$
IQV $13.5$ $3.69$ $8.29$ $4.04$
CNV $6.98$ $2.83$ $5.87$ $3.41$
Table 6: Uncertainty statistics for AF classifier
The WVR, SDM and CNV give equivalent values as this is a binary classification problem, and their expressions are equivalent in this case (see Appendix B). However, we observe a similar pattern to the uncertainty statistics computed for the land cover case study. Namely that the entropy is the most sensitive to variations in the predictive probability distributions, with UVR being the least sensitive.
4.3 Summary of results
The application of our ML classification uncertainty framework across different models and datasets allowed evaluation of the predictive PMFs, whilst taking into account the various sources of uncertainty described in Figure 1. All uncertainty statistics of nominal properties described in Section 3.2.3 were evaluated based on the outputted predictive PMFs, with results given in Tables 45, and Table 6.
The results show consistent behaviour in the predictive PMFs, in that more accurate ML classification models produced uncertainty statistics with values close to zero. The results also showed that the information entropy was the most sensitive overall to variations in the PMF out of all uncertainty statistics considered, with UVR being the most robust. However, if one is more concerned with assessing highly uncertain PMFs, the IQV is the most sensitive. We also demonstrated that three uncertainty statistics (WVR, SDM, and CNV) are equivalent in the case of binary classification, and approximately equivalent for multi-class ML classification models with highly confident predictive PMFs.
@@ 5 Conclusion
In this paper, we have presented a metrological framework for evaluating uncertainty of nominal properties, motivated by the increasing use of ML classification models in a wide range of applications. We have identified and listed a number of sources of uncertainty within ML classification models which contribute to the output predictive uncertainty of the nominal property or variable, namely training data uncertainty, class assignment uncertainty, uncertainty about choice of model, model parameter uncertainty, and input data uncertainty, only some of which are currently addressed in the GUM suite of documents.
The work described in this paper has focused primarily on predictive uncertainty of ML classification models, with a specific focus on probabilistic ML classification models, i.e. models that return class probabilities for the output nominal variable. This focus included the evaluation of uncertainty for such nominal variables, which is not currently considered within the VIM and GUM. We have demonstrated that the PMF provides a complete characterisation of the uncertainty associated with nominal properties, and thus the predictive uncertainty in probabilistic ML classifiers. We have also motivated the use of the PMF of a nominal variable, as opposed to a summary statistic, in propagating such uncertainty through a multistage measurement model, representing either a metrological traceability chain or a processing chain for information products. We then provided example approaches using analytic or sampling methods of how to perform such propagation using the PMF.
We have reviewed various statistics that have been proposed within the literature associated with expressing the predictive uncertainty of probabilistic ML classifiers. We provided a comparison between these statistics in terms of their fundamental properties, and applied them in the context of two case studies that rely upon a metrological analysis, namely land cover classification, and atrial fibrillation detection. We found that for both case studies, the information entropy, which is the most common statistic for uncertainty within the ML community, is in general, the most sensitive to variations in the predictive probability distributions. We have also shown that three of the uncertainty statistics (WVR, SDM and CNV) are equivalent for binary classification, and approximately equivalent in the multi-class case for confident ML classification model outputs. We have also illustrated how sources of uncertainty are addressed within the two case studies using different methodologies dependent on the classification modelling procedure.
The framework we have outlined builds upon and enhances previous work in uncertainty evaluation for nominal properties. The framework enables an extension of the GUM to uncertainty evaluation for nominal properties, and provides methods for handling the sources of uncertainty within ML classification, underpinning their use within metrological applications.
@@ Acknowledgements
The original data used for LC classification study can be found at [82, 83, 84, 85]. More details on how this data was processed for the LC classification study are provided in [17]. The original data for the case study on AF detection (DeepBeat) can be found at [86] and is described in [7]. The custom data split used was described in [75].
Appendix A Derivation of WVR and UVR in terms of expected distance from mode
We can express the WVR and UVR defined in Section 3.2.3 in terms of a simple distance metric from the mode(s), defined as
$d_{k}=\begin{cases}0&p_{k}=\widehat{p},\\ 1&\text{otherwise},\end{cases}\quad k=1,\dots,K.$ (A.1)
Wilcox’s Variation Ratio
Using the fact that $\widehat{p}=1-\mathbf{d}\cdot\mathbf{p}$ for unimodal distributions, the expression for WVR (3.12) becomes
$u_{\text{WVR}}(\mathbf{p})$ $=1-\frac{K\widehat{p}-1}{K-1}$
$=1-\frac{K(1-\mathbf{d}\cdot\mathbf{p})-1}{K-1}$
$=\frac{K}{K-1}\mathbf{d}\cdot\mathbf{p}$ (A.2)
Since the probability vector $\mathbf{p}$ represents a PMF, we see that $\mathbf{d}\cdot\mathbf{p}=\mathbb{E}_{\mathbf{p}}[\mathbf{d}]$. This is the expected distance from the mode. Thus, we can express the normalised WVR in terms of the expected distance
$u_{\text{WVR}}(\mathbf{p})=\frac{K}{K-1}\mathbb{E}_{\mathbf{p}}[\mathbf{d}]$ (A.3)
We see a direct analogy with the variance for continuous variables in terms of the expected squared distance from the mean: $\text{Var}[X]=\mathbb{E}_{p}[d(X)^{2}]$, where $d(X)=X-\mathbb{E}_{p}[X]$.
Universal Variation Ratio
For multimodal distributions, $m\widehat{p}=1-\mathbb{E}_{\mathbf{p}}[\mathbf{d}]$, where $m$ is the number of modes. In this case, the UVR in expression (3.14) can be written as
$u_{\text{UVR}}(\mathbf{p})$ $=\frac{K^{2}}{K^{2}-1}\left(1-\frac{\widehat{p}}{m}\right)$
$=\frac{K^{2}}{K^{2}-1}\left(1-\frac{1-\mathbb{E}_{\mathbf{p}}[\mathbf{d}]}{m^{2}}\right)$
$=\frac{K^{2}}{(K^{2}-1)m^{2}}\mathbb{E}_{\mathbf{p}}[\mathbf{d}]+\frac{K^{2}}{K^{2}-1}\frac{m^{2}-1}{m^{2}}$ (A.4)
Thus the UVR can be expressed as a linear transformation of the expected distance from the mode even when the PMF is multi-modal, whereas expression (A.3) is only valid for unimodal distributions.
Appendix B Relationship between WVR, SDM and CNV
Although expressions for the uncertainty statistics WVR (3.12), SDM (3.16) and CNV (3.25) give different values in general, there is a relationship between them for certain forms of PMF $\mathbf{p}$.
Consider the unimodal case where the modal probability is $\widehat{p}$, and all other class probabilities are equal. This PMF can be written as
$q_{k}=\begin{cases}\widehat{p}&c_{k}=\widehat{c},\\ \frac{1-\widehat{p}}{K-1}&\text{otherwise},\end{cases}\quad k=1,\dots K.$ (B.1)
All binary class ($K=2$) PMFs can be written in this form. For multi-class classification models where outputs are confident in one class, the predictive output PMFs approximately follow the form of (B.1).
We will now show that WVR, SDM and CNV uncertainty statistics are equivalent in such a scenario. Substituting (B.1) into (3.16) leads to
$u_{\text{SDM}}(\mathbf{q})$ $=1-\sqrt{\dfrac{(K-1)(\widehat{p}-\frac{1-\widehat{p}}{K-1})^{2}}{K-1}}$
$=1-\widehat{p}+\frac{1-\widehat{p}}{K-1}$
$=1-\frac{K\widehat{p}-1}{K-1}=u_{\text{WVR}}(\mathbf{q}).$
Now substituting (B.1) into (3.25) yields
$u_{\text{CNV}}(\mathbf{q})$ $=1-\sqrt{\frac{K\left(\widehat{p}^{2}+(K-1)\left(\frac{1-\widehat{p}}{K-1}\right)^{2}\right)-1}{K-1}}$
$=1-\frac{1}{K-1}\sqrt{(K-1)(K\widehat{p}^{2}-1)+K(1-\widehat{p})^{2}}$
$=1-\frac{1}{K-1}\sqrt{K^{2}\widehat{p}^{2}-2K\widehat{p}+1}$
$=1-\frac{1}{K-1}\sqrt{(K\widehat{p}-1)^{2}}=u_{\text{WVR}}(\mathbf{q})$
This also explains the equivalence in the binary class uncertainty expressions (3.13), (3.17), and (3.27).